# Minimmit: Fast Finality with Even Faster Blocks

### v4

@@ 1. Introduction
1.1. Evidence for reduced latency
@@ 1.3. Dropping the slow path
Kudzu. Kudzu makes the more general assumption that $n\geq 3f+2p+1$ for a tunable parameter $p$. Of course, in the case that $p=f$, this corresponds to the same assumption that $n\geq 5f+1$. The protocol guarantees liveness and consistency so long as at most $f$ processors display Byzantine faults. It also guarantees that, during synchrony, a correct leader can finalise a block after a single round of voting, so long as the number of faulty processors is $\leq p$. For a correct leader during synchrony, the ‘slow path’ (requiring two rounds of voting) is therefore only relevant if the number of faulty processors is strictly between $p$ and $f$. In this case, the quorum required for finality after a single round of voting is also larger than for Minimmit (npn-p rather than nfn-f). Kudzu therefore gives no strict ‘like-for-like’ improvement in resilience compared to Minimmit, although the tunable parameter that Kudzu provides is beneficial.
@@ 3. The intuition
One round of voting. We aim to specify a standard form of view-based protocol, in which each view has a designated leader. If a view has a correct leader and begins after GST, the leader should send a block $b$ to all other processors, who will then send a signed vote for $b$ to all others. Upon receipt of nfn-f votes for $b$ (by distinct processors), our intention is that processors should then be able to immediately finalise $b$: this is called 2-round finality (Martin and Alvisi, 2006) (one round to send the block and one round for voting). $5f-1\leq n$ is necessary and sufficient [^6] See (Kuznetsov et al., 2021) and https://decentralizedthoughts.github.io/2021-03-03-2-round-bft-smr-with-n-equals-4-f-equals-1/ for 2-round finality: we make the assumption that $5f+1\leq n$ for simplicity. We’ll call a set of nfn-f votes for a block an L-notarisation (where ‘L’ stands for ‘large’).
When to enter the next view? We specified above that a single L-notarisation should suffice for finality. However, as well as achieving 2-round finality, we also wish processors to proceed to the next view immediately upon seeing a smaller set of votes, which we’ll call an M-notarisation (where the ‘M’ stands for ‘mini’). As explained in Section 1, the rationale behind this is that, in realistic scenarios, receiving nfn-f votes will generally take much longer than receiving a significantly smaller set of votes (of size $2f+1$, say). Proceeding to view $v+1$ immediately upon seeing an M-notarisation for $b$ in view $v$ therefore speeds up the process of block formation: processors can begin view $v+1$ earlier, and then finalise $b$ upon later receiving the larger notarisation.
@@
Nullifications. Since it may be the case that no block for view $v$ receives an M-notarisation (e,g., if the leader is Byzantine), processors must sometimes produce signed messages indicating that they wish to move to view $v+1$ because of a lack of progress in view $v$: we’ll call these nullify(v)(v) messages. [^8] Our approach here is somewhat similar to Simplex (Chan and Pass, 2023), but the reader need not have familiarity with that protocol to follow the discussion. We do not yet specify precisely when processors send $\text{nullify}(v)$ messages (we will do so shortly). For now, we promise only that a statement analogous to (X1) will hold for $\text{nullify}(v)$ messages:
@@ 4. The formal specification
7:
8:   If $\mathtt{S}$ contains a valid proposal $b$ for view $\mathtt{v}$: $\triangleright$ As defined in Section 4
9:    If $\mathtt{notarised}=\bot$ and $\mathtt{nullified}=$ false:
10:    Set $\mathtt{notarised}:=b$ and send $(\text{vote},b)$ to all processors; $\triangleright$ Send vote
11:
12:   If $\mathtt{T}=2\Delta$, $\mathtt{nullified}=$ false and $\mathtt{notarised}=\bot$:
13:    Set $\mathtt{nullified}:=$true and send $(\text{nullify},\mathtt{v})$ to all processors; $\triangleright$ Send nullify(v)(v)
14:
15:   If $\mathtt{S}$ contains a nullification for $\mathtt{v}$:
16:    Set $\mathtt{v}:=\mathtt{v}+1$, $\mathtt{nullified}:=$ false, $\mathtt{proposed}=$ false, $\mathtt{notarised}:=\bot$; $\triangleright$ Go to next view
17:   If $\mathtt{S}$ contains an M-notarisation for some $b$ with $b.\text{view}=\mathtt{v}$:
18:    If $\mathtt{notarised}=\bot$ and $\mathtt{nullified}=$ false, send $(\text{vote},b)$ to all processors; $\triangleright$ Send vote
19:    Set $\mathtt{v}:=\mathtt{v}+1$, $\mathtt{nullified}:=$ false, $\mathtt{proposed}=$ false, $\mathtt{notarised}:=\bot$; $\triangleright$ Go to next view
20:
21:   If $\mathtt{nullified}=$ false, $\mathtt{notarised}\neq\bot$ and $\mathtt{S}$ contains $\geq 2f+1$ messages, each signed by a
22:   different processor, and each either:
23:   (i) A message $(\text{nullify},\mathtt{v})$, or;
24:   (ii) Of the form $(\text{vote},b)$ for some $b$ s.t. $b.\text{view}=\mathtt{v}$ and $\mathtt{notarised}\neq b$:
25:   Set $\mathtt{nullified}:=$ true and send $(\text{nullify},\mathtt{v})$ to all processors;
@@
27:
28:   If $\mathtt{S}$ contains a new L-notarisation for any block $b$:
29:    Finalise $b$; $\triangleright$ Finalisation as specified in Section 2
@@ 5.1. Consistency
Recall that $\bot$ is a default value, different than any block. Each correct processor’s local value $\mathtt{notarised}$ is initially set to $\bot$, and is also set to $\bot$ upon entering any view (lines 16 and 19). A correct processor $p_{i}$ will only vote for a block if $\mathtt{notarised}=\bot$ (lines 9 and 18). The claim of the lemma holds because, upon voting for any block $b$, $p_{i}$ either sets $\mathtt{notarised}:=b$ (line 10) and then does not redefine this value until entering the next view, or else immediately enters the next view (lines 18 and 19). ∎
@@
Towards a contradiction, suppose $b$ receives an L-notarisation, $v=b.\text{view}$, and view $v$ receives a nullification. Let $P$ be the correct processors that vote for $b$, let $P^{\prime}=\Pi\setminus P$, and note that $|P^{\prime}|\leq 2f$. Since view $v$ receives a nullification, it follows that some processor in $P$ must send a nullify(v)(v) message. So, let $t$ be the first timeslot at which some processor $p_{i}\in P$ sends such a message. Since $p_{i}$ cannot send a nullify(v)(v) message upon timeout (lines 12-13), $p_{i}$ must send the nullify(v)(v) message at $t$ because the conditions of lines 21-24 hold for $p_{i}$ at $t$, i.e., $p_{i}$ must have received $\geq 2f+1$ messages, each signed by a different processor, and each of the form:
@@
From Lemma 5.2, it follows that $v_{2}>v_{1}$. According to clause (ii) from the definition of when $\mathtt{S}$ contains a valid proposal for view $v_{2}$, correct processors will not vote for $b_{2}$ in line 10 until they receive an M-notarisation for its parent, $b_{0}$ say. Correct processors will not vote for $b_{2}$ in line 18 until $b_{2}$ has already received an M-notarisation, meaning that at least $f+1$ correct processors must first vote for $b_{2}$ via line 10, and $b_{0}$ must receive an M-notarisation. By our choice of $v_{2}$, it follows that $b_{0}.\text{view}<v_{1}$. This gives a contradiction, because, by clause (iii) from the definition of a valid proposal for view $v_{2}$, correct processors would not vote for $b_{2}$ in line 10 without receiving a nullification for view $v_{1}$. By Lemma 5.3, such a nullification cannot exist. So, block $b_{2}$ cannot receive an M-notarisation (and no correct processor votes for $b_{2}$ via either line 10 or 18). ∎
@@ 5.2. Liveness
So, the conditions of lines 21-24 are eventually satisfied, meaning that $p_{j}$ sends a nullify(v)(v) message (line 25). Any correct processor that does not vote for a view $v$ block also sends a nullify(v)(v) message, so all correct processors send nullify(v)(v) messages. All correct processors therefore receive a nullification for view $v$ and leave the view (line 16), giving the required contradiction. ∎
@@
Suppose $p_{i}=\mathtt{lead}(v)$ is correct and that the first correct processor $p_{j}$ to enter view $v$ does so at timeslot $t\geq\text{GST}$. If $v>1$, processor $p_{j}$ enters view $v$ upon receiving either a nullification for view v1v-1, or else an M-notarisation for some view v1v-1 block. Since $p_{j}$ forwards on all new notarisations and nullifications that it receives (lines 2 and 3), it follows that all correct processors enter view $v$ by $t+\Delta$ (note that this also holds if $v=1$). Processor $p_{i}$ therefore sends a new block $b$ to all processors by $t+\Delta$, which is received by all processors by $t+2\Delta$. Let $b^{\prime}$ be the parent of $b$ and suppose $b^{\prime}.\text{view}=v^{\prime}$. Then $p_{i}$ receives an M-notarisation for $b^{\prime}$ by $t+\Delta$. Since $p_{i}$ forwards on all new notarisations that it receives (line 3), all correct processors receive an M-notarisation for $b^{\prime}$ by $t+2\Delta$. Since $p_{i}$ has entered view $v$, it must also have received nullifications for all views in the open interval $(v^{\prime},v)$ by $t+\Delta$, and all correct processors receive these by $t+2\Delta$. All correct processors therefore vote for $b$ (by either line 10 or 18) [^12] The point of line 18 is to ensure this part of the proof goes through. As noted in Section 3, without it, there is the possibility that correct processors move to the next view upon seeing an M-notarisation, before they are able to vote via line 10, thereby failing to guarantee an L-notarisation. before any correct processor sends a nullify(v)(v) message. The block $b$ therefore receives an L-notarisation, as claimed. ∎
@@ 5.3. Optimistic responsiveness
As before, suppose first $p_{i}=\mathtt{lead}(v)$ is correct and that the first correct processor to enter view $v$ does so at timeslot $t\geq\text{GST}$. Since correct processors forward on all new notarisations and nullifications that they receive, it follows that all correct processors enter view $v$ by $t+\delta$. Processor $p_{i}$ therefore sends a new block $b$ to all processors by $t+\delta$, which is received by all processors by $t+2\delta$. Let $b^{\prime}$ be the parent of $b$ and suppose $b^{\prime}.\text{view}=v^{\prime}$. Then, from the definition of the function SelectParent, it follows that $p_{i}$ receives an M-notarisation for $b^{\prime}$ by $t+\delta$. Since $p_{i}$ forwards on all new notarisations that it receives (line 3), all correct processors receive an M-notarisation for $b^{\prime}$ by $t+2\delta$. Since $p_{i}$ has entered view $v$, it must also have received nullifications for all views in the open interval $(v^{\prime},v)$ by $t+\delta$, and all correct processors receive these by $t+2\delta$. All correct processors therefore vote for $b$ (by either line 10 or 18) by $t+2\delta$, and before any correct processor sends a nullify(v)(v) message. All correct processors therefore receive $b$ together with an L-notarisation (and an M-notarisation) for $b$ by $t+3\delta$, and also leave view $v$ by this time. This establishes the claim of the lemma. ∎
@@
So, the conditions of lines 21-24 are satisfied at this time, meaning that $p_{j}$ sends a nullify(v)(v) message (line 25). Any correct processor that does not vote for a view $v$ block also sends a nullify(v)(v) message by this time. So, all correct processors receive a nullification for view $v$ by $t+2\Delta+3\delta$, giving the required contradiction. ∎
@@ 8. Related work
Modern 2-Round Protocols. Alpenglow (Kniep et al., 2025) is formally analysed under the assumption that $n\geq 5f+1$. The paper also considers circumstances in which the protocol can tolerate a further $f$ crash failures, but the required assumptions for this case (essentially that Byzantine leaders cannot carry out a form of proposal equivocation) do not hold under partial synchrony. Banyan (Vonlanthen et al., 2024) carries out the fast path in parallel with the slow path mechanism, but can suffer from unbounded message complexity with faulty leaders. Kudzu (like FaB) makes the more general assumption that $n\geq 3f+2p+1$ for a tunable parameter $p$. Kudzu and Alpenglow also describe the use of erasure coding techniques to allow for significantly improved maximum throughput (such techniques were also employed by SMR protocols in earlier papers, such as (Shoup, 2023)).

### v5

@@ 1. Introduction
$1.$ Evidence for reduced latency
@@ 1.3. Dropping the slow path
Kudzu. Kudzu makes the more general assumption that $n\geq 3f+2p+1$ for a tunable parameter pp. Of course, in the case that $p=f$, this corresponds to the same assumption that $n\geq 5f+1$. The protocol guarantees liveness and consistency so long as at most $f$ processors display Byzantine faults. It also guarantees that, during synchrony, a correct leader can finalise a block after a single round of voting, so long as the number of faulty processors is $\leq p$. For a correct leader during synchrony, the ‘slow path’ (requiring two rounds of voting) is therefore only relevant if the number of faulty processors is strictly between $p$ and $f$. In this case, the quorum required for finality after a single round of voting is also larger than for Minimmit (npn-p rather than nfn-f). Kudzu therefore gives no strict ‘like-for-like’ improvement in resilience compared to Minimmit, although the tunable parameter that Kudzu provides is beneficial.
@@ 3. The intuition
One round of voting. We aim to specify a standard form of view-based protocol, in which each view has a designated leader. If a view has a correct leader and begins after GST, the leader should send a block $b$ to all other processors, who will then send a signed vote for $b$ to all others. Upon receipt of nfn-f votes for $b$ (by distinct processors), our intention is that processors should then be able to immediately finalise $b$: this is called 2-round finality (Martin and Alvisi, 2006) (one round to send the block and one round for voting). $5f-1\leq n$ is necessary and sufficient [^6] See (Kuznetsov et al., 2021) and https://decentralizedthoughts.github.io/2021-03-03-2-round-bft-smr-with-n-equals-4-f-equals-1/ for 2-round finality: we make the assumption that $5f+1\leq n$ for simplicity. We’$l$ call a set of nfn-f votes for a block an L-notarisation (where ‘L’ stands for ‘large’).
When to enter the next view? We specified above that a single L-notarisation should suffice for finality. However, as well as achieving 2-round finality, we also wish processors to proceed to the next view immediately upon seeing a smaller set of votes, which we’$l$ call an M-notarisation (where the ‘M’ stands for ‘mini’). As explained in Section 1, the rationale behind this is that, in realistic scenarios, receiving nfn-f votes will generally take much longer than receiving a significantly smaller set of votes (of size $2f+1$, say). Proceeding to view $v+1$ immediately upon seeing an M-notarisation for $b$ in view $v$ therefore speeds up the process of block formation: processors can begin view $v+1$ earlier, and then finalise $b$ upon later receiving the larger notarisation.
@@
Nullifications. Since it may be the case that no block for view $v$ receives an M-notarisation (e,g., if the leader is Byzantine), processors must sometimes produce signed messages indicating that they wish to move to view $v+1$ because of a lack of progress in view $v$: we’$l$ call these nullify(v)(v) messages. [^8] Our approach here is somewhat similar to Simplex (Chan and Pass, 2023), but the reader need not have familiarity with that protocol to follow the discussion. We do not yet specify precisely when processors send $\text{nullify}(v)$ messages (we will do so shortly). For now, we promise only that a statement analogous to (X1) will hold for $\text{nullify}(v)$ messages:
@@ 4. The formal specification
7:    Set $\mathtt{proposed}:=$ true;
8:
9:   If $\mathtt{S}$ contains a valid proposal $b$ for view $\mathtt{v}$: $\triangleright$ As defined in Section 4
10:    If $\mathtt{notarised}=\bot$ and $\mathtt{nullified}=$ false:
11:    Set $\mathtt{notarised}:=b$ and send $(\text{vote},b)$ to all processors; $\triangleright$ Send vote
12:
13:   If $\mathtt{T}=2\Delta$, $\mathtt{nullified}=$ false and $\mathtt{notarised}=\bot$:
14:    Set $\mathtt{nullified}:=$true and send $(\text{nullify},\mathtt{v})$ to all processors; $\triangleright$ Send nullify(v)(v)
15:
16:   If $\mathtt{S}$ contains a nullification for $\mathtt{v}$:
17:    Set $\mathtt{v}:=\mathtt{v}+1$, $\mathtt{nullified}:=$ false, $\mathtt{proposed}:=$ false, $\mathtt{notarised}:=\bot$, $\mathtt{T}:=0$;
@@
19:   If $\mathtt{S}$ contains an M-notarisation for some $b$ with $b.\text{view}=\mathtt{v}$:
20:    If $\mathtt{notarised}=\bot$ and $\mathtt{nullified}=$ false, send $(\text{vote},b)$ to all processors; $\triangleright$ Send vote
21:    Set $\mathtt{v}:=\mathtt{v}+1$, $\mathtt{nullified}:=$ false, $\mathtt{proposed}:=$ false, $\mathtt{notarised}:=\bot$, $\mathtt{T}:=0$;
[…]
23:
24:   If $\mathtt{nullified}=$ false, $\mathtt{notarised}\neq\bot$ and $\mathtt{S}$ contains $\geq 2f+1$ messages, each signed by a
25:   different processor, and each either:
26:   (i) A message $(\text{nullify},\mathtt{v})$, or;
27:   (ii) Of the form $(\text{vote},b)$ for some $b$ s.t. $b.\text{view}=\mathtt{v}$ and $\mathtt{notarised}\neq b$:
28:   Set $\mathtt{nullified}:=$ true and send $(\text{nullify},\mathtt{v})$ to all processors;
[…]
30:
31:   If $\mathtt{S}$ contains a new L-notarisation for any block $b$:
32:    Finalise $b$; $\triangleright$ Finalisation as specified in Section 2
@@ 5.1. Consistency
Recall that $\bot$ is a default value, different than any block. Each correct processor’s local value $\mathtt{notarised}$ is initially set to $\bot$, and is also set to $\bot$ upon entering any view (lines 17 and 21). A correct processor $p_{i}$ will only vote for a block if $\mathtt{notarised}=\bot$ (lines 10 and 20). The claim of the lemma holds because, upon voting for any block $b$, $p_{i}$ either sets $\mathtt{notarised}:=b$ (line 11) and then does not redefine this value until entering the next view, or else immediately enters the next view (lines 20 and 21). ∎
@@
Towards a contradiction, suppose $b$ receives an L-notarisation, $v=b.\text{view}$, and view $v$ receives a nullification. Let $P$ be the correct processors that vote for $b$, let $P^{\prime}=\Pi\setminus P$, and note that $|P^{\prime}|\leq 2f$. Since view $v$ receives a nullification, it follows that some processor in $P$ must send a nullify(v)(v) message. So, let $t$ be the first timeslot at which some processor $p_{i}\in P$ sends such a message. Since $p_{i}$ cannot send a nullify(v)(v) message upon timeout (lines 13-14), $p_{i}$ must send the nullify(v)(v) message at $t$ because the conditions of lines 24-27 hold for $p_{i}$ at $t$, i.e., $p_{i}$ must have received $\geq 2f+1$ messages, each signed by a different processor, and each of the form:
@@
From Lemma 5.2, it follows that $v_{2}>v_{1}$. According to clause (ii) from the definition of when $\mathtt{S}$ contains a valid proposal for view $v_{2}$, correct processors will not vote for $b_{2}$ in line 11 until they receive an M-notarisation for its parent, $b_{0}$ say. Correct processors will not vote for $b_{2}$ in line 20 until $b_{2}$ has already received an M-notarisation, meaning that at least $f+1$ correct processors must first vote for $b_{2}$ via line 11, and $b_{0}$ must receive an M-notarisation. By our choice of $v_{2}$, it follows that $b_{0}.\text{view}<v_{1}$. This gives a contradiction, because, by clause (iii) from the definition of a valid proposal for view $v_{2}$, correct processors would not vote for $b_{2}$ in line 11 without receiving a nullification for view $v_{1}$. By Lemma 5.3, such a nullification cannot exist. So, block $b_{2}$ cannot receive an M-notarisation (and no correct processor votes for $b_{2}$ via either line 11 or 20). ∎
@@ 5.2. Liveness
So, the conditions of lines 24-27 are eventually satisfied, meaning that $p_{j}$ sends a nullify(v)(v) message (line 28). Any correct processor that does not vote for a view $v$ block also sends a nullify(v)(v) message, so all correct processors send nullify(v)(v) messages. All correct processors therefore receive a nullification for view $v$ and leave the view (line 17), giving the required contradiction. ∎
@@
Suppose $p_{i}=\mathtt{lead}(v)$ is correct and that the first correct processor $p_{j}$ to enter view $v$ does so at timeslot $t\geq\text{GST}$. If $v>1$, processor $p_{j}$ enters view $v$ upon receiving either a nullification for view v1v-1, or else an M-notarisation for some view v1v-1 block. Since $p_{j}$ forwards on all new notarisations and nullifications that it receives (lines 2 and 3), it follows that all correct processors enter view $v$ by $t+\Delta$ (note that this also holds if $v=1$). Processor $p_{i}$ therefore sends a new block $b$ to all processors by $t+\Delta$, which is received by all processors by $t+2\Delta$. Let $b^{\prime}$ be the parent of $b$ and suppose $b^{\prime}.\text{view}=v^{\prime}$. Then $p_{i}$ receives an M-notarisation for $b^{\prime}$ by $t+\Delta$. Since $p_{i}$ forwards on all new notarisations that it receives (line 3), all correct processors receive an M-notarisation for $b^{\prime}$ by $t+2\Delta$. Since $p_{i}$ has entered view $v$, it must also have received nullifications for all views in the open interval $(v^{\prime},v)$ by $t+\Delta$, and all correct processors receive these by $t+2\Delta$. All correct processors therefore vote for $b$ (by either line 11 or 20) [^12] The point of line 20 is to ensure this part of the proof goes through. As noted in Section 3, without it, there is the possibility that correct processors move to the next view upon seeing an M-notarisation, before they are able to vote via line 11, thereby failing to guarantee an L-notarisation. before any correct processor sends a nullify(v)(v) message. The block $b$ therefore receives an L-notarisation, as claimed. ∎
@@ 5.3. Optimistic responsiveness
As before, suppose first $p_{i}=\mathtt{lead}(v)$ is correct and that the first correct processor to enter view $v$ does so at timeslot $t\geq\text{GST}$. Since correct processors forward on all new notarisations and nullifications that they receive, it follows that all correct processors enter view $v$ by $t+\delta$. Processor $p_{i}$ therefore sends a new block $b$ to all processors by $t+\delta$, which is received by all processors by $t+2\delta$. Let $b^{\prime}$ be the parent of $b$ and suppose $b^{\prime}.\text{view}=v^{\prime}$. Then, from the definition of the function SelectParent, it follows that $p_{i}$ receives an M-notarisation for $b^{\prime}$ by $t+\delta$. Since $p_{i}$ forwards on all new notarisations that it receives (line 3), all correct processors receive an M-notarisation for $b^{\prime}$ by $t+2\delta$. Since $p_{i}$ has entered view $v$, it must also have received nullifications for all views in the open interval $(v^{\prime},v)$ by $t+\delta$, and all correct processors receive these by $t+2\delta$. All correct processors therefore vote for $b$ (by either line 11 or 20) by $t+2\delta$, and before any correct processor sends a nullify(v)(v) message. All correct processors therefore receive $b$ together with an L-notarisation (and an M-notarisation) for $b$ by $t+3\delta$, and also leave view $v$ by this time. This establishes the claim of the lemma. ∎
@@
So, the conditions of lines 24-27 are satisfied at this time, meaning that $p_{j}$ sends a nullify(v)(v) message (line 28). Any correct processor that does not vote for a view $v$ block also sends a nullify(v)(v) message by this time. So, all correct processors receive a nullification for view $v$ by $t+2\Delta+3\delta$, giving the required contradiction. ∎
@@ 8. Related work
Modern 2-Round Protocols. Alpenglow (Kniep et al., 2025) is formally analysed under the assumption that $n\geq 5f+1$. The paper also considers circumstances in which the protocol can tolerate a further $f$ crash failures, but the required assumptions for this case (essentially that Byzantine leaders cannot carry out a form of proposal equivocation) do not hold under partial synchrony. Banyan (Vonlanthen et al., 2024) carries out the fast path in parallel with the slow path mechanism, but can suffer from unbounded message complexity with faulty leaders. Kudzu (like FaB) makes the more general assumption that $n\geq 3f+2p+1$ for a tunable parameter pp. Kudzu and Alpenglow also describe the use of erasure coding techniques to allow for significantly improved maximum throughput (such techniques were also employed by SMR protocols in earlier papers, such as (Shoup, 2023)).